SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Links
Meetings
Accelerator Systems Division
Accelerator Operations Reports
MCC Log Book
PEP Physics Log Book
Hypernews
Babar/PEP-II Liason
Background Remediation
Beam Size Measurement
Beam-Beam
Luminosity
SVT Radiation
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

PEP News from MCC

Here is a short and non-exhaustive summary of the PEP meetings held on Tuesday 03 and Thursday 05 April at MCC.

Tuesday Meeting :

  • LER x-tune has been lowered, closed to 1/2 integer (last thursday).
  • HER LFB still not working.
  • HER BPM8012 not working... blind HER Y-angle.
  • hints of vacuum burst in the HER.
  • Plans:
  • fix HER LFB/BPM
  • Try to run HER RF 12-6
  • Install PM skew quads (gain 30% LER vertical emittance)
  • continue the fast dither commissioning
  • increase HER RF voltage

PEP down: leak in LCLS construction area!

  • Tuesday: all water lines (LCW and domestic) turned off: all magnets down: Pictures
  • Wednesday: leak found on domestic pipe: fixed in the evening.
  • Thursday: magnets turned on.
  • Beams on: